OpenStudy (anonymous):

Could someone please double check these answers quick? Thanks! :) 1. Evaluate lim (as t approaches infinity) sin(1/t) Answer: 0 2. The motion of an object along a horizontal path is modeled by the equation x(t) = 3 - |t - 3|, where x(t) is the position from the origin in meters and t is the time in seconds. Find the distance traveled by the object on the time interval 0 ≤ t ≤ 10. Answer: 4 3. Locate the absolute maximum value of the function f (x) = x^3 - 27x + k on the interval [0, 5]. Answer: k-10

OpenStudy (astrophysics):

1. Looks good

OpenStudy (anonymous):

Yay! :)

OpenStudy (astrophysics):

May I ask how you approached 2? Your answer does look correct

OpenStudy (astrophysics):

x(10) right?

OpenStudy (astrophysics):

You also did realize that distance is a scalar quantity

OpenStudy (anonymous):

Yep! :)

OpenStudy (astrophysics):

So 2. Looks good

OpenStudy (anonymous):

d = x(10) = (3 - |10 - 3|)m = (3 - |7|)m = -4m

OpenStudy (astrophysics):

Right but its +4 as distance is a scalar quantity

OpenStudy (anonymous):

Yepperz :)

OpenStudy (astrophysics):

How did you get k-10? What does that mean?

OpenStudy (astrophysics):

What is the first derivative of f(x)?

OpenStudy (anonymous):

f'(x) = 3x² - 27 = 0

OpenStudy (astrophysics):

Right now solve for x and see what values are within your given interval then we must take the second derivative that will tell us if it's a maximum or minimum

OpenStudy (astrophysics):

(second derivative at the x value)

OpenStudy (anonymous):

3x² = 27 x² = 27/3 = 9 x = ±√9 = ±3 I used x=3 f"(x) = 6x f"(3) = 6(3) = 18 f(0) = (0)³ - 27(0) + k = k f(5) = (5)³ - 27(5) + k = -10 + k Thus, the maximum occurs at x = 0 but my options given are: k-54 k-10 k there is no absolute maximum value so I went with k-10 since that's what I got before reaching 0

OpenStudy (zarkon):

I don't agree with your answer to #2

OpenStudy (anonymous):

What do you think it is?? :)

OpenStudy (zarkon):

no

OpenStudy (zarkon):

they are not asking for the distance from the start. They are asking for the distance traveled

OpenStudy (anonymous):

My options for #4 are: 3 4 7 10

OpenStudy (zarkon):

I know the answer...I don't need a list

OpenStudy (anonymous):

Okay sorry... Just thought that might help

OpenStudy (anonymous):

How would I find the distance traveled?

OpenStudy (zarkon):



OpenStudy (astrophysics):

Ah gotcha, sorry I was thinking of where it started

OpenStudy (astrophysics):

@Breee615 What is the distance travelled used Zarkon's image if needed

OpenStudy (anonymous):

3?? :)

OpenStudy (anonymous):

*10 gosh I can't type today

OpenStudy (anonymous):

2. The motion of an object along a horizontal path is modeled by the equation x(t) = 3 - |t - 3|, where x(t) is the position from the origin in meters and t is the time in seconds. Find the distance traveled by the object on the time interval 0 ≤ t ≤ 10. (Distance is different from displacement) x'(t)=(t-3)/|3-t| (from the def. of abs value) x'(t) is negative for t>3, so if you just go, x(10)-x(0)? then this is not precise, because you are looking for scalar (not vector) displacement. Rather, |x(0)-x(3)| + |x(3)-x(10)| is the distance, i think

OpenStudy (anonymous):

oh, my bad wrong f

OpenStudy (zarkon):

yes...10 is a better answer

OpenStudy (anonymous):

Okay that makes a bit more sense, thank you :)

OpenStudy (anonymous):

x'(t)=(t-3)/|t-3| (for the correct f) but, then x'(t)<0 for x<3, so you have the same distance, |x(0)-x(3)| + |x(3)-x(10)|

OpenStudy (anonymous):

((just that I misread the function))

OpenStudy (anonymous):

Does my answer for 3 make sense? :)

OpenStudy (anonymous):

@astrophysics

OpenStudy (anonymous):

@fakeee